OpenStudy (anonymous):

The Perimeter of a rectangle is 52. Its length is 2 more than twice the measure of its width. find its dimensions.

OpenStudy (anonymous):

Equation: 52 = 2(2w + 2) + 2w

OpenStudy (anonymous):

Distribute the 2 outside the parentheses to both terms INSIDE the parentheses. Afterwards combine like terms. Solve for w. What ever you get for w, plug it into the equation below for w: L = 2(w) + 2

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

No problem. Take care

OpenStudy (anonymous):

perimeter ofrectangle=2(l+b) as l=2b per=2(3b)=52 6b=52 b=52/6 b=8.6 l=2b l=2(8.6) l-17.8

OpenStudy (anonymous):

length=17.8 breadth=8.67